OpenStudy (caroline1218):

The ordered pairs (1, 1), (2, 8), (3, 27), (4, 64), and (5, 125) represent a function. What is a rule that represents this function?

OpenStudy (caroline1218):

A. y = x B. y = 4x C. y = x^2 D. y = x^3

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

try the conditions one by one. A says that both numbers in a bracket are equal. is this the case?

OpenStudy (caroline1218):

No, both numbers are not equal for each pair.

OpenStudy (anonymous):

B says this: first number times 4 = second number what about this one?

OpenStudy (anonymous):

then C says that: first times first = second and D says that: first times first times first = second so check each possibility

OpenStudy (caroline1218):

So it's D?

OpenStudy (anonymous):

@caroline1218 yes

OpenStudy (caroline1218):

Thank you :)

OpenStudy (anonymous):

you're welcome :)

OpenStudy (anonymous):

The ordered pairs contain x and y-values. (2,8) means x = 2, y = 8